OpenStudy (anonymous):

Solve 4e^(3x) + 2 = 18. 1/3ln4 ln e3 ln 12 ln 7

OpenStudy (freckles):

Have you tried to isolate the exponential part?

OpenStudy (freckles):

that is first subtract 2 on both sides then divide both sides by 4

OpenStudy (anonymous):

Okay one sec

OpenStudy (anonymous):

I have e^(3x) = 4

OpenStudy (dan815):

yep good

OpenStudy (freckles):

ok and knowing ln(e^u)=u take ln( ) of both sides

OpenStudy (dan815):

now can you rewrite 4 as a power of 3

OpenStudy (dan815):

power of e* sorry lol

OpenStudy (dan815):

that was dyslexia on another level

OpenStudy (anonymous):

so it would be e^4x= 3?

OpenStudy (dan815):

\[e^{3x}=4\] okay u are good upto here right

OpenStudy (anonymous):

yes

OpenStudy (dan815):

now \[4= e^{something}\]

OpenStudy (dan815):

you can just use the ln function but im trying to make you understand how bases work

OpenStudy (dan815):

\[e^{3x}=4\\ ln e^{3x}= ln 4\\ 3x= ln4\\ x=ln(4)/3\]

OpenStudy (anonymous):

Ohh okay so adding the ln to both sides removed the e

OpenStudy (dan815):

ln function is the inverse of e

OpenStudy (dan815):

do you know how logarithms work

OpenStudy (anonymous):

Only vaguely

OpenStudy (dan815):

\[ln(y)=x ===> y=e^x\] means x is the exponent on e to get y

OpenStudy (anonymous):

Okay that makes sense. Thanks a ton!

OpenStudy (anonymous):

So just making sure its A?